# Beyond Syntax: Action Semantics Learning for App Agents

### v1

@@ Abstract
The advent of Large Language Models (LLMs) enables the rise of App agents that interpret user intent and operate smartphone Apps through actions such as clicking and scrolling. While prompt-based solutions with closed LLM APIs show promising ability, they incur heavy compute costs and external API dependency. Fine-tuning smaller open-source LLMs solves these limitations. However, current fine-tuning methods use a syntax learning paradigm that forces agents to reproduce exactly the ground truth action strings, leading to out-of-distribution (OOD) vulnerability. To fill this gap, we propose Action Semantics Learning (ASL), a novel learning framework, where the learning objective is capturing the semantics of the ground truth actions. Specifically, inspired by the programming language theory, we define the action semantics for App agents as the state transition induced by the action in the user interface. With this insight, ASL employs a novel SEmantic Estimator (SEE) to compute a semantic reward to train the App agents in generating actions aligned with the semantics of ground truth actions, even when the syntactic forms differ. To support the effectiveness of ASL, we theoretically demonstrate the superior robustness of ASL for the OOD problem compared with the existing syntax learning paradigm. Extensive experiments on offline and online smartphone App operation benchmarks show that ASL significantly improves the accuracy and generalisation of App agents over existing methods.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs)  [^1] In this work, LLMs refer to foundation models that process multiple input modalities (e.g., visual or multimodal) and generate textual outputs. has led to the emergence of AI agents across a wide range of domains li2023camel ; goutora ; ramos2025review ; hu2025selfevolving . Among these, AI agents for smartphone App operation, commonly referred to as App agents, are gaining significant attention for leveraging LLMs to enable fully automated App navigation rawles2024androidinthewild ; rawles2024androidworld ; mobileagentv2 ; wang2025distrl ; chen2025spa ; christianos2025lightweight . These agents interpret user-provided goals in natural language and execute tasks such as scheduling, messaging, and online purchasing by interacting with the user interfaces (UIs) via human-like actions, such as tapping, scrolling, or entering text. The diversity of smartphone operating systems and UI designs, coupled with predefined action sets for specific systems, poses a significant challenge in developing generalisable App agents.
@@
Recent advances in App agents can be broadly grouped into two approaches. One prominent approach leverages closed LLMs APIs, such as GPT and Gemini, through prompt engineering techniques to comprehend user intents and device states, subsequently generating suitable actions rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . While these App agents exhibit robust reasoning and generalisation capabilities, their practical deployment is often hindered by challenges such as high inference latency, substantial computational resource requirements, and frequent dependence on third-party APIs, which may escalate operational costs and introduce instability. Alternatively, fine-tuning smaller LLMs with offline datasets or interactive online trajectories has emerged as a viable strategy, achieving low-latency inference conducive to real-world applications wang2025distrl ; christianos2025lightweight ; bai2024digirl . However, current methods typically follow a syntax learning paradigm that forces agents to reproduce exactly the ground truth action string for each UI within the training data and ignores the underlying action semantics. This leads to an out-of-distribution (OOD) vulnerability: even the UIs slightly differ from the samples within the training data can provoke a dramatic performance drop. An agent trained to follow a specific interface path may fail when the UI changes or alternative options are presented. For example, an agent trained to delete text by generating "Tap the Delete button" will fail if the Delete button is unavailable, whilst the "Cut" button is still visible. See Sec. 3.1 for further discussion of this OOD vulnerability.
Figure 1: Our action semantics learning framework.
@@
To go beyond syntax and capture action semantics, we first draw inspiration from the denotational semantics in programming language theory, which defines program semantics by its effect on system state stoy1981denotational , to define the action semantics for App agents as the state transition induced by the action in the UI environment. Building on this insight, we propose Action Semantics Learning (ASL), a novel learning framework that shifts the learning objective of App agents from exactly reproducing the ground truth action strings to reproducing their semantics. Specifically, during training, the actions generated by the App agent are rewarded by a SEmantic Estimator (SEE) that scores how closely the semantics of the generated action match that of the ground truth action. With this semantic reward, ASL allocates credit to all actions yielding the same UI state transition as the ground truth action, teaching the agent to prioritise action semantics over action strings. In Sec. 3.4, we theoretically demonstrate that, compared with the syntax learning paradigm, ASL is more robust towards the OOD problem. See Fig. 1 for a visualisation of the proposed ASL framework.
@@
The semantic estimator SEE is the core component of our learning framework. Specifically, SEE leverages an LLM-driven world model to predict the impact of an action on the current UI state and uses BERT to measure the semantic similarity between the effects of the predicted action and the ground truth action. Our design is inspired by recent world-model–empowered web agents web-world , which leverage predicted next observations as offline rewards for prompt engineering. In contrast, we are the first to incorporate these world-model predictions directly into the app agent training stage, generating a semantic reward without requiring any extra computational resources for deployment.
@@
$\bullet$ We conduct extensive experiments on both offline and online smartphone operation benchmarks. Empirical results demonstrate that the proposed ASL framework achieves superior accuracy and generalisation compared to existing approaches.
@@ 2.1 Prompt-driven App Agents
Prompt-driven App agents mark a substantial advancement in human–device interaction, enabling natural language commands to be translated into precise UI actions on smartphones, tablets, and other mobile devices. Recent research has progressively refined these systems, with pioneering implementations like AppAgent yang2023appagent demonstrating that large pretrained models can effectively map screenshots or accessibility trees to human-like actions. Building on these foundations, more recent frameworks have introduced advanced prompting strategies incorporating planning, self-reflection, and tool use to tackle increasingly complex tasks rawles2024androidworld ; mobileagentv2 . Despite their strong performance, current prompt-driven App agents remain heavily dependent on closed LLM APIs such as GPT and Gemini rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . This dependence introduces several limitations, including high inference latency, substantial computational costs, and reliance on third-party services that raise concerns around data privacy, service reliability, and vendor lock-in. These challenges hinder real-world deployment and highlight the need for open, efficient, and controllable alternatives.
2.2 Fine-tuned App Agents
@@
Recent research has increasingly focused on fine-tuning smaller models tailored for mobile environments. LiMAC (christianos2025lightweight, ) introduces a compact action transformer that predicts UI-grounded actions paired with a fine-tuned vision–language model for textual reasoning, while InfiGUIAgent (infiguiagent, ) adopts a two-stage approach that first learns UI semantics from screenshots and then generates actions conditioned on user instructions. Despite architectural differences, these methods follow a syntax learning paradigm that forces agents to reproduce the exact ground truth action string. This paradigm can lead to OOD vulnerability: even slight UI differences from the training data samples can provoke dramatic performance degradation. Complementing these offline approaches, several works explore online optimisation to improve generalisability and robustness to OOD data. DigiRL (bai2024digirl, ) proposes a reinforcement learning framework that simulates App control scenarios and fine-tunes policies from an offline-supervised starting point. DistRL (wang2025distrl, ) enhances this approach with asynchronous online training, boosting sample efficiency. However, these online approaches often require extensive interaction with real mobile device environments during training, significantly increasing computational and operational costs. For instance, DigiRL’s framework involves building a scalable and parallelisable Android learning environment equipped with a vision-language model-based evaluator, necessitating substantial resources for real-device emulators. Similarly, DistRL employs centralised training with decentralised data acquisition, requiring multiple device interactions to collect diverse trajectories. In our work, we aim to enhance model generalisability without requiring interaction with real device environments during training.
2.3 World Model boosted App Agents
@@
World models simulate future states of the environment based on previous states and given actions lecun2022path ; ding2024understanding . In domains such as robotics, world models yang2023learning ; zhen20243d ; zhou2024robodreamer ; ajay2023compositional can predict future visual states, such as images or videos across diverse environments. These predictive capabilities are critical for enabling robots to understand their surroundings, make informed decisions, and accurately execute tasks. Building on these advances, researchers have extended world modelling to web environments, where models web-world ; isyour ; liu2023picture generate textual predictions of a website’s future state following a user interaction, enabling the development of more accurate web agents. However, existing applications of world models in this setting are typically used for offline feedback, which incurs additional inference overhead and is suboptimal for practical deployment. In our work, we instead leverage world models as a reward signal to fine-tune the agent by computing the semantic similarity between predicted and actual actions, requiring no additional inference time or computational resources during deployment.
@@ 3 Methodology
In this section, we begin with the problem formulation, which introduces the smartphone app operation task and summarises the key challenge for current fine-tuning methods. Then, we propose a novel learning framework named Action Semantic Learning (ASL) for App agents. Furthermore, we elaborate on the semantic estimator, the key component within the ASL framework. Finally, we conduct a theoretical analysis to support the effectiveness of the ASL framework.
@@ 3.1 Problem Formulation
Smartphone operation. We formulate smartphone operation as a Partially Observable Markov Decision Process defined by $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O},\gamma)$, where $\mathcal{S}$ represents smartphone system states, $\mathcal{A}$ denotes actions for operating UIs, $\mathcal{T}$ is the state transition function, $\mathcal{R}$ is the utility function, $\mathcal{O}$ maps states to UI observations, and $\gamma$ is the discount factor. At each timestep $t$, the App agent, which is parameterised by $\theta$ and denoted as $\pi_{\theta}$, receives an observation $o_{t}$ of the current UI state, the user goal $g$ in natural language, and action history $h_{t}$. The App agent produces an action according to $a^{\theta}_{t}=\pi_{\theta}(g,h_{t},o_{t})$, where the observation $o_{t}$ may include a subset of available screenshots and textual information like UI element trees android_control . Let $\mathcal{D}_{train}=\{(g_{i},h_{i},o_{i},a_{i}^{*})\}_{i=1}^{N}$ denote a training dataset with $N$ step samples, where $a_{i}^{*}$ is the ground truth action for sample $i$. Our objective is to fine-tune an App agent to maximise the success rate of completing the user goal $g$, while also considering generalisation and robustness. For simplicity, we assume $\gamma=1$. The utility function $\mathcal{R}$ returns 1 when the user goal can be completed within the required maximum action steps, and 0 otherwise.
OOD vulnerability for syntax learning. Current approaches for training App agents typically employ a syntax learning paradigm, where models learn to reproduce exactly the ground truth action string rawles2024androidworld ; mobileagentv2 ; yang2023appagent ; android_control . Mathematically, their learning objective can be formulated as:
@@
where $\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\log p_{\theta}(a_{i}^{*}\mid g_{i},h_{i},o_{i})$ is the supervised fine-tuning (SFT) loss function for the data sample $i$ and $p_{\theta}(a_{i}^{*}|g_{i},h_{i},o_{i})$ is the likelihood of the ground truth action under the App agent policy $\pi_{\theta}$ for the given inputs. The key limitation of this paradigm is that it rewards the model only when the output exactly matches the ground truth action, while penalising all other outputs, even those that share identical semantics with the ground truth action. Consequently, the model optimises for token-level cross-entropy rather than understanding action semantics, leading to severe out-of-distribution (OOD) vulnerability: even slight UI differences from the training samples can cause performance drops. For example, Table 3 in Sec. 4 shows that on the AndroidControl benchmark android_control , the accuracy of the App agent fine-tuned with Llama-3 plunges from 65.5% on the in-distribution dataset (IDD) to just 57.1% on the App-Unseen OOD set. Mathematically, we formulate this OOD vulnerability as Theorem 3.1 with the proof in the Appendix.
@@
Let $\pi^{\mathrm{SFT}}_{\theta}$ be an App agent trained under the syntax-learning objective of Eq. (1); for an $n$-step task with ground truth actions $\{a_{1}^{*},\dots,a_{n}^{*}\}$, define $P(\text{success})=\Pr[\text{the agent completes all $n$ steps correctly}]$ and, for any semantically preserving but syntactically non-trivial perturbation $\delta$ satisfying $\delta(a_{i}^{*})\neq a_{i}^{*}$ for at least one $i$, define $P(\text{success}\mid\delta)=\Pr[\text{the agent completes all $n$ steps correctly when each }a_{i}^{*}\text{ is replaced by }\delta(a_{i}^{*})]$; letting $\Delta(\delta)=P(\text{success})-P(\text{success}\mid\delta)$, we have $\Delta(\delta)>0$.
@@ 3.2 Action Semantics Learning
Action semantics. Inspired by the denotational semantics in programming language theory, which defines program semantics by its effect on system state stoy1981denotational , we mathematically define the action semantics for App agents as the following:
@@
$J(\theta)=\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big{[}\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\big{]}\Big{]},$ (2)
@@
To optimise the objective function in Eq. (2), we leverage the REINFORCE algorithm (williams1992simple, ). This yields the following gradient estimator:
$\nabla_{\theta}J(\theta)\;=\;\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big{[}\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\,\nabla_{\theta}\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big{]}\Big{]},$
@@
$\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big{[}r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\,\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big{]}.$ (3)
@@
Semantics-aware fine-tuning. Optimising the semantics-aware loss function in Eq. (3) enables App agents to learn action semantics by rewarding semantically equivalent actions. However, computing this loss requires sampling actions from the agent and scoring these actions using a reward model, which can lead to prohibitively high training costs.
To maintain efficiency, inspired by previous works in the literature nakano2021webgpt ; ouyang2022training ; rafailov2023direct , we limit the number of sampled actions as one per data sample. However, sampling only a single action can lead to vanishing gradients when the sampled action is not semantically equivalent to the ground truth, hindering effective learning and potentially leading to sub-optimal performance. To address this, we propose a combined loss function that integrates our semantics-aware loss function with the SFT loss function. Concretely, for each data sample, the loss function is formulated as follows:
$\ell^{ASL}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})+\lambda_{i}\,\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*}),$ (4)
@@
Intuitively, the combined loss function operates as follows: (1) when the sampled action is semantically correct but syntactically different, learning is primarily guided by the semantic-aware loss; (2) when the action matches both semantics and syntax, both loss components contribute jointly; and (3) when the action matches neither, the SFT loss dominates, ensuring stable gradient flow even in the absence of semantic reward. This design effectively mitigates gradient vanishing issues while preserving training efficiency, as it requires sampling only a single action from the App agent per training instance. We provide a more detailed discussion for this intuition in the Appendix.
In the next section, we propose a novel SEmantic Estimator (SEE) for App agents, which serves as the reward model in our ASL framework.
3.3 Semantic Estimator
During offline training, the direct interaction with a real smartphone UI environment is often impractical. Therefore, it is challenging to get accurate semantic rewards, namely the probability that an agent-generated action is semantically equivalent to the ground truth action. To address this limitation, we introduce a SEmantic Estimator (SEE) that consists of two key components: (1) a world model that simulates the smartphone UI state transitions, and (2) a semantic calculator that estimates the semantic reward by quantifying the similarity between the transitions.
World model. Inspired by recent advances in world-model–based web agents web-world , which leverage predicted future observations as rewards for prompt engineering, SEE utilises a world model to mimic the state transition function $\mathcal{T}(s,a)$ mentioned in Definition 3.1. Mathematically, the workflow of this world model can be formulated as:

### v2

@@ Abstract
The recent development of Large Language Models (LLMs) enables the rise of App agents that interpret user intent and operate smartphone Apps through actions such as clicking and scrolling. While prompt-based solutions with proprietary LLM APIs show promising ability, they incur heavy compute costs and external API dependency. Fine-tuning smaller open-source LLMs solves these limitations. However, current supervised fine-tuning methods use a syntax learning paradigm that forces agents to reproduce exactly the ground truth action strings, leading to out-of-distribution (OOD) vulnerability. To fill this gap, we propose Action Semantics Learning (ASL), a novel learning framework, where the learning objective is capturing the semantics of the ground truth actions. Specifically, inspired by the programming language theory, we define the action semantics for App agents as the state transition induced by the action in the user interface. Building on this insight, ASL employs a novel SEmantic Estimator (SEE) to compute a semantic similarity to train the App agents in generating actions aligned with the semantics of ground truth actions, even when their syntactic forms differ. SEE is a flexible module that can be applied in both supervised and reinforcement fine-tuning paradigms. To support the effectiveness of ASL, we theoretically demonstrate the superior robustness of ASL for the OOD problem compared with the existing syntax learning paradigm. Extensive experiments across multiple offline and online benchmarks demonstrate that ASL significantly improves the accuracy and generalisation of App agents compared to existing methods.
@@ 1 Introduction
The rapid development of Large Language Models (LLMs)  [^1] In this work, LLMs refer to foundation models that process multiple input modalities (e.g., visual or multimodal) and generate textual outputs. has led to the emergence of AI agents across a wide range of domains [24]  [15, 9]. Among these, AI agents for smartphone App operation, commonly referred to as App agents, are gaining significant attention for leveraging LLMs to enable fully automated App navigation [28, 27, 30, 4]. These agents interpret user-provided goals in natural language and execute tasks such as scheduling, messaging, and online purchasing by interacting with the user interfaces (UIs) via pre-defined actions, such as tapping, scrolling, or entering text. The diversity of smartphone operating systems and UI descriptions, coupled with predefined action sets for specific systems, poses a significant challenge in developing generalisable App agents.
@@
Recent advances in App agents can be broadly grouped into two approaches. One prominent approach leverages proprietary LLMs APIs, such as GPT [13] and Gemini [11], through prompt engineering techniques to comprehend user intents and device states, subsequently generating suitable actions [27, 16]. While these App agents exhibit robust reasoning and generalisation capabilities, their practical deployment is often hindered by challenges such as high inference latency, substantial computational resource requirements, and frequent dependence on third-party APIs, which may escalate operational costs and introduce instability. Alternatively, fine-tuning smaller LLMs with offline datasets or interactive online trajectories has emerged as a viable strategy, achieving low-latency inference conducive to real-world applications [31, 5, 2]. However, current supervised methods typically follow a syntax learning paradigm that forces agents to reproduce exactly the ground truth action string for each UI within the training data, instead of understanding the underlying action semantics. This leads to an out-of-distribution (OOD) vulnerability: even the UIs slightly differ from the samples within the training data can provoke a dramatic performance drop. An agent trained to follow a specific interface path may fail when the UI changes or alternative options are presented. For example, an agent trained to delete text by generating "Tap the Delete button" will fail if the Delete button is unavailable, whilst the "Cut" button is still visible. See Sec. 3.1 for further discussion of this OOD vulnerability.
Figure 1: Our ASL framework. The reward corresponds to the function defined in Eq. (2) .
@@
To go beyond syntax and capture action semantics, we first draw inspiration from the denotational semantics in programming language theory, which defines program semantics by its effect on system state [29], to define the action semantics for App agents as the state transition induced by the action in the UI environment. Building on this insight, we propose Action Semantics Learning (ASL), a novel learning framework that shifts the learning objective of App agents from exactly reproducing the ground truth action strings to reproducing their semantics. Specifically, during training, the actions generated by the App agent are rewarded by a SEmantic Estimator (SEE) that scores how closely the semantics of the generated action match that of the ground truth action. By employing the semantic reward, ASL considers actions that lead to the same UI state transition as the ground-truth action to be semantically equivalent, thereby guiding the agent to focus on semantics rather than syntactic form. In Sec. 3.4, we theoretically demonstrate that, compared with the syntax learning paradigm, ASL is more robust towards the OOD problem. See Fig. 1 for a visualisation of the proposed ASL framework.
@@
The semantic estimator SEE is the core component of our learning framework. Specifically, SEE leverages an LLM-driven world model to predict the impact of an action on the current UI state and uses BERT to measure the semantic similarity between the effects of the predicted action and the ground truth action. Our design is inspired by recent world‐model–empowered web agents [3], which leverage predicted observations as offline rewards for prompt engineering. In contrast, we are the first to incorporate world‐model predictions directly into the app agent training stage, generating a semantic reward without requiring any extra computational resources for deployment.
@@
$\bullet$ We conduct extensive experiments on various online and offline benchmarks. Empirical results demonstrate that the proposed ASL framework achieves superior accuracy and generalisation compared to existing approaches, especially on more difficult tasks.
@@ 2 Related Works
Prompt-driven App agents. Prompt-driven App agents mark a substantial advancement in human–device interaction, enabling natural language commands to be translated into precise UI actions on smartphones, tablets, and other mobile devices. Recent research has progressively refined these systems, with pioneering implementations like AppAgent [39] demonstrating that large pretrained models can effectively map screenshots or accessibility trees to human-like actions. Building on these foundations, more recent frameworks have introduced advanced prompting strategies incorporating planning, self-reflection, and tool use to tackle increasingly complex tasks [27, 30, 32]. Despite their strong performance, current prompt-driven App agents remain heavily dependent on proprietary LLM APIs such as GPT and Gemini [27, 30, 39, 16]. This dependence introduces several limitations, including high inference latency, substantial computational costs, and reliance on third-party services that raise concerns around data privacy, service reliability, and vendor lock-in. These challenges hinder real-world deployment and highlight the need for open, efficient, and controllable alternatives. Unlike these prompt-driven agents, our approach directly learns action semantics during fine-tuning, reducing reliance on inference-time prompting and enabling more robust and lightweight execution.
@@
Fine-tuned App agents. Recent research has focused on fine-tuning smaller models tailored for mobile environments. LiMAC [5] introduces a compact action transformer that predicts UI-grounded actions paired with a fine-tuned vision–language model for textual reasoning, while InfiGUIAgent [18] adopts a two-stage approach that first learns UI semantics from screenshots and then generates actions conditioned on user instructions. Despite architectural differences, these supervised fine-tuning methods follow a syntax learning paradigm that forces agents to reproduce the exact ground truth action string. This paradigm can lead to OOD vulnerability: even slight UI differences from the training data samples can provoke dramatic performance degradation. Complementing these offline approaches, several works explore online optimisation to improve generalisability and robustness to OOD data. DigiRL [2] proposes a reinforcement learning framework that simulates App control scenarios and fine-tunes policies from an offline-supervised starting point. DistRL [31] enhances this approach with asynchronous online training, boosting sample efficiency. However, such online methods typically assign rewards based on sparse binary outcomes from the environment, without assessing individual steps. In contrast, our SEE module introduces step-level evaluation, providing finer-grained and smoother feedback at each step rather than relying exclusively on environment-level rewards.
@@
World model boosted App agents. World models simulate future states of the environment based on previous states and given actions [14, 8]. In domains such as robotics, world models [37, 42, 44, 1] can predict future visual states, such as images or videos across diverse environments. These predictive capabilities are critical for enabling robots to understand their surroundings, make informed decisions, and accurately execute tasks. Building on these advances, researchers have extended world modelling to web environments, where models [3, 10, 17] generate textual predictions of a website’s future state following a user interaction, enabling the development of more accurate web agents. However, existing applications of world models in this setting are typically utilised for providing feedback at test time, which introduces additional inference overhead and remains suboptimal for practical deployment. In our work, we instead leverage world models as a reward signal to fine-tune the agent by computing the semantic similarity between predicted and ground truth actions, requiring no additional inference time or computational resources in deployment.
@@ 3 Methodology
In this section, we begin with the problem formulation, which introduces the smartphone app operation task and summarises the key challenge for current fine-tuning methods. Then, we propose a novel learning framework named Action Semantic Learning (ASL) for App agents. Furthermore, we elaborate on the semantic estimator (SEE), the key component within the ASL framework. Finally, we conduct a theoretical analysis to support the effectiveness of the ASL framework.
@@ 3.1 Problem Formulation
Smartphone operation. We formulate smartphone operation as a Partially Observable Markov Decision Process defined by $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\mathcal{O},\gamma)$, where $\mathcal{S}$ represents smartphone system states, $\mathcal{A}$ denotes actions for operating UIs, $\mathcal{T}$ is the state transition function, $\mathcal{R}$ is the utility function, $\mathcal{O}$ maps states to UI observations, and $\gamma$ is the discount factor. At each timestep $t$, the App agent, which is parameterised by $\theta$ and denoted as $\pi_{\theta}$, receives an observation $o_{t}$ of the current UI state, the user goal $g$ in natural language, and action history $h_{t}$. The App agent produces an action according to $a^{\theta}_{t}=\pi_{\theta}(g,h_{t},o_{t})$, where the observation $o_{t}$ may include a subset of available screenshots and textual information like UI element trees [16]. Let $\mathcal{D}_{train}=\{(g_{i},h_{i},o_{i},a_{i}^{*})\}_{i=1}^{N}$ denote a training dataset with $N$ step samples, where $a_{i}^{*}$ is the ground truth action for sample $i$. Our objective is to fine-tune an App agent to maximise the success rate of completing the user goal $g$, while also considering generalisation and robustness. For simplicity, we assume $\gamma=1$. The utility function $\mathcal{R}$ returns 1 when the user goal can be completed in the required maximum action steps, and 0 otherwise.
OOD vulnerability for syntax learning. Current supervised approaches for training App agents typically employ a syntax learning paradigm, where models learn to reproduce exactly the ground truth action string [27, 30, 39, 16]. Mathematically, their learning objective can be formulated as:
@@
where $\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\log p_{\theta}(a_{i}^{*}\mid g_{i},h_{i},o_{i})$ is the supervised fine-tuning (SFT) loss function for the data sample $i$ and $p_{\theta}(a_{i}^{*}|g_{i},h_{i},o_{i})$ is the likelihood of the ground truth action under the App agent policy $\pi_{\theta}$ for the given inputs. The key limitation of this paradigm is that it rewards the model only when the output exactly matches the ground truth action, while penalising all other outputs, even those that share identical semantics with the ground truth action. Consequently, the model optimises for token-level cross-entropy rather than understanding action semantics, leading to severe out-of-distribution (OOD) vulnerability: even minor and naturally occurring variations in UIs can be treated as OOD, causing the model to overfit to training patterns and degrade in generalisation. We formulate this OOD vulnerability as Theorem 3.1 with the proof in the supplementary material.
@@
[…]
@@ 3.2 Action Semantics Learning
Action semantics. Inspired by the denotational semantics in programming language theory, which defines program semantics by its effect on system state [29], we mathematically define the action semantics for App agents as the following:
@@
$J(\theta)=\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big[$ (2)
$\quad\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big[r(a,a^{\star}_{i},g_{i},h_{i},o_{i})\big]\Big],$
@@
To optimise the objective function in Eq. (2), we leverage the REINFORCE algorithm [33]. This yields the following gradient estimator:
$\nabla_{\theta}J(\theta)\;=\;\mathbb{E}_{(g_{i},h_{i},o_{i},a_{i}^{*})\sim\mathcal{D}_{train}}\Big[$ (3)
$\quad\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big[r(a,a^{\star}_{i},g_{i},h_{i},o_{i})$
$\quad\cdot\nabla_{\theta}\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big]\Big],$
@@
$\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=-\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid g_{i},h_{i},o_{i})}\big[r(a,a^{\star}_{i},g_{i},h_{i},o_{i})$ (4)
$\quad\cdot\log p_{\theta}(a\mid g_{i},h_{i},o_{i})\big].$
@@
Supervised fine-tuning with SEE. Optimising the semantics-aware loss function in Eq. (4) enables App agents to learn action semantics by rewarding semantically equivalent actions. However, computing this loss requires sampling actions from the agent and scoring these actions using a reward model, which can lead to prohibitively high training costs.
To maintain efficiency, inspired by previous works in the literature [20, 21, 26], we limit the number of sampled actions as one per data sample. However, sampling only a single action can lead to vanishing gradients when the sampled action is not semantically equivalent to the ground truth, hindering effective learning and potentially leading to sub-optimal performance. To address this, we propose a combined loss function that integrates our semantics-aware loss function with the SFT loss function. The loss function is formulated as follows:
$\ell^{ASL}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})=\ell^{S}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*})$ (5)
$\quad+\lambda_{i}\,\ell^{SFT}_{\theta}(g_{i},h_{i},o_{i},a_{i}^{*}),$
@@
Intuitively, the combined loss function operates as follows: (1) when the sampled action is semantically correct but syntactically different, learning is primarily guided by the semantic-aware loss; (2) when the action matches both semantics and syntax, both loss components contribute jointly; and (3) when the action matches neither, the SFT loss dominates, ensuring stable gradient flow even in the absence of semantic reward. This design efficiently mitigates gradient vanishing issues, as it requires sampling only a single action from the App agent per training instance.
Reinforcement fine-tuning with SEE. Beyond supervised fine-tuning, we follow DigiRL [2] and optimise the policy with advantage-weighted regression (AWR) [23]. For each transition $(g_{i},h_{i},o_{i},a_{i})$, we inject a thresholded semantic reward $\tilde{r}(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i})=\mathds{1}\{r(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i})>\tau\}\,r(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i})$, where $\tilde{a}_{i}$ is the teacher model-predicted reference action and $\tau$ is a semantic threshold. We then define the augmented reward $r^{\mathrm{see}}(\cdot)$ as:
$r^{\mathrm{see}}(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i})=r^{\mathrm{env}}(a_{i},g_{i},h_{i},o_{i})$ (6)
$\quad+\beta\,\tilde{r}(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i}),$
[…]
$A^{\mathrm{DigiRL\text{-}SEE}}(g_{i},h_{i},o_{i},a_{i},\tilde{a}_{i})=\eta^{H-i}\,r^{\mathrm{env}}(a_{H},g_{H},h_{H},o_{H})$ (7)
$\quad+\big(1-\eta^{H-i}\big)(V^{\mathrm{step}}(g_{i+1},h_{i+1},o_{i+1})$
$\quad+r^{\mathrm{see}}(a_{i},\tilde{a}_{i},g_{i},h_{i},o_{i})-V^{\mathrm{step}}(g_{i},h_{i},o_{i})),$
where $H$ is the horizon, $\eta\!\in\![0,1]$ balances the high-variance Monte Carlo estimate $\eta^{H-i}\,r^{\mathrm{env}}(a_{H},g_{H},h_{H},o_{H})$ and the value-based baseline from a learned doubly-rubost estimator $V^{\mathrm{step}}$. The optimisation objective of the policy $\pi_{\theta}$ is a hard-filtered AWR loss as:
$\begin{aligned} &\ell^{\mathrm{DigiRL\text{-}SEE}}_{\theta}(g_{i},h_{i},o_{i},a_{i},\tilde{a}_{i})=-\,\mathds{1}\!\left\{A^{\mathrm{DigiRL\text{-}SEE}}(g_{i},h_{i},o_{i},a_{i},\tilde{a}_{i})>0\right\}\\ &\quad\cdot\log p_{\theta}(a_{i}\mid g_{i},h_{i},o_{i}).\end{aligned}$
SEE augments DigiRL with step-level semantic feedback, thereby improving credit assignment under sparse, binary outcomes and yielding stronger generalisation on long-horizon UI control. In the next section, we elaborate on the Semantic Estimator (SEE) for App agents.
3.3 SEE: Semantic Estimator
To define the semantics of an action, we focus on its effect rather than its string syntax or associated metadata. Predefined actions often come in varying formats, and typically include parameters such as the ID of a UI element that lacks standalone semantic meaning, especially without the associated user input or context. In contrast, we propose that the result of executing an action within a given state carries rich semantic information, and that this outcome can serve as a more meaningful definition of the action itself. We introduce a SEmantic Estimator (SEE) that consists of two key components: (1) a world model that simulates the smartphone UI state transitions, and (2) a semantic calculator that estimates the semantic reward by quantifying the similarity between the transitions.
World model. Inspired by recent advances in world-model–based web agents [3], which leverage predicted future observations as rewards for prompt engineering, SEE utilises a world model to mimic the state transition function $\mathcal{T}(s,a)$ mentioned in Definition 3.1. Mathematically, the workflow of this world model can be formulated as: